# Survey on Factuality in Large Language Models: Knowledge, Retrieval and Domain-Specificity

### v1

@@ Abstract
Index Terms:
Large language Models, Factuality, Knowledge, Retrieval, Domain-Specificity, Evaluation, Enhancement
1 Introduction
[…]
@@ 1 Introduction
, leaf, text width=22em ] ] [ Retrieval Aug-
mented LLMs [ BingChat (Microsoft, 2023), Llama-Index (Liu, 2022), , leaf, text width=22em ] ] ] [ Impact (Sec 2.3) [ Legal [ Curran et al. (2023), Pranshu Verma (2023), Sands (2023) , leaf, text width=22em ] ] [ Medicine [ Nori et al. (2023), Shen et al. (2023), Kung et al. (2023) , leaf, text width=22em ] ] ] ] [ Evaluation
@@
C-Eval (Huang et al., 2023b), AGIEval (Zhong et al., 2023a),
RealTimeQA (Kasai et al., 2022), FreshQA (Vu et al., 2023), leaf, text width=33em ] ] [ Factuality
@@
Wang et al. (2023a), Yu et al. (2023), Kadavath et al. (2022), leaf, text width=33em ] ] [ Domain-Factuality
Evaluation (Sec 3.4) [ FLARE (Finance) (Xie et al., 2023b), CMB (Medicine) (Wang et al., 2023d),
GeneTuring (Gene) (Jin et al., 2023), EcomInstruct (Ecom) (Li et al., 2023e), leaf, text width=33em ] ] ] [ Analysis
@@
Immemorization, Forgetting, Reasoning Failure, leaf, text width=24em ] ] [ Retrieval-Level 4.2.2 [ Insufficient Information, Distracting Information,
@@
Misinterpretation of Related Information, leaf, text width=24em ] ] [ Inference-Level 4.2.3 [ Snowballing, Erroneous Decoding, Exposure Bias, leaf, text width=24em ] ] ] ] [ Enhancement
@@
, leaf, text width=24em ] ] [ SFT-based 5.1.2 [ Continual SFT: Moiseev et al. (2022);
@@
, leaf, text width=22em ] ] [ Multi-Agent 5.1.3 [ Multi-debate: Du et al. (2023)
@@
Agent-based Retrieval: ReAct (Yao et al., 2023a) , leaf, text width=24em ] ] [ Retrieval
@@
, leaf, text width=24em ] ] [ Retrieval on External
Memory 5.2.4 [ General Knowledge: Li et al. (2022c),
@@
Entity Knowledge: knowledge plugin (Zhang et al., 2023g)
@@
, leaf, text width=24em ] ] [ Retrieval on
Structured Data 5.2.5 [ StructGPT (Jiang et al., 2023a) , Zhang et al. (2023d) , leaf, text width=24em ] ] ] [ Domain Factuality
@@
Finance: EcomGPT (Li et al., 2023e), BloombergGPT (Wu et al., 2023)
@@
, leaf, text width=30em ] ] ] ]
Figure 1: Taxonomy of research in factuality in Large Language Models (LLMs) that consists of the issue, causes, evaluation, analysis, and enhancement methods.
@@
Amidst these pioneering efforts, the emergence of Large Language Models (LLMs), such as ChatGPT (OpenAI, 2022b), GPT-4 (OpenAI, 2023) and LLaMA (Touvron et al., 2023a, b), has been seen as a significant leap in both academics and industries, especially towards AI systems possessing vast factual knowledge (OpenAI, 2022a, 2023). The advantages of using LLMs as knowledge bases carriers are manifold. Firstly, they reduce the overhead and costs associated with building and maintaining dedicated knowledge bases (Petroni et al., 2019c; AlKhamissi et al., 2022; Wang et al., 2023b). Additionally, LLMs offer a more flexible approach to knowledge processing and utilization, allowing for context-aware reasoning and the ability to adapt to novel information or prompts (Sun et al., 2023a; Huang and Chang, 2023). Yet, with their unparalleled capabilities, concerns have arisen about the potential of LLMs to generate non-factual or misleading content (Bender et al., 2021; OpenAI, 2023; Bubeck et al., 2023). In light of these advancements and challenges, this survey seeks to delve deeply into the LLMs, exploring both their potential and the concerns surrounding their factual accuracy.
@@
Understanding the factuality of Large Language Models is more than just a technical challenge; it’s essential for the responsible use of these tools in our daily lives. As LLMs become more integrated into services like search engines (Microsoft, 2023), chatbots (OpenAI, 2022b; Google, 2023), and content generators (Cui et al., 2023b), the information they provide directly influences decisions, beliefs, and actions of millions of people. If an LLM provides incorrect or misleading information, it can lead to misunderstandings, spread false beliefs, or even cause harm, especially for those domains that demand high factual accuracy (Ling et al., 2023b), such as health  (Thirunavukarasu et al., 2023; Tang et al., 2023), law (Huang et al., 2023a), and finance (Wu et al., 2023). For instance, a physician relying on an LLM for medical guidance might inadvertently jeopardize patient health, a corporation leveraging LLM insights might make ill-informed market decisions, or an attorney misinformed by an LLM might falter in legal proceedings (Curran et al., 2023). In addition, with the advancement of LLM-based agents, the factuality of LLMs is becoming even more potent. A driver or an autonomous driving car might rely on LLM-based agents for planning or driving, where serious factual mistakes made by LLMs could cause irreversible damage. By studying the factuality of LLMs, we aim to ensure that these models are both powerful and trustworthy.
@@
A surge of research has been directed towards evaluating LLMs’ factuality, which encompasses diverse tasks like factoid question answering and fact checking. Beyond evaluation, efforts to improve the factual knowledge of LLMs have been notable. Strategies have ranged from retrieving information from external knowledge bases to continual pretraining and supervised finetuning. Yet, despite these burgeoning efforts, a holistic overview that covers the full spectrum of factuality in LLMs remains elusive. While there are existing surveys in the field, such as those by Chang et al. (2023) and Wang et al. (2023g), that delve into the evaluation of LLMs and their factuality, they only scratch the surface of the broader landscape. There are also a bunch of recent studies focusing on hallucinations in LLMs (Rawte et al., 2023; Zhang et al., 2023f; Rawte et al., 2023; Ye et al., 2023). But we differentiate between the hallucination issue and the factuality issues in Sec 2.2. Moreover, these surveys often overlook key areas we emphasize, like domain-specific factuality or the challenge of outdated information. While Ling et al. (2023a) explores domain specialization in LLMs, our survey takes a more expansive look at the broader issues of factuality. To the best of our understanding, our work is the first comprehensive study on the factuality of large language models.
@@
This survey aims to offer an exhaustive overview of the factuality studies in LLMs, delving into four key dimensions: Sec 2) The definition and impact of the factuality issue (Pranshu Verma, 2023; Nori et al., 2023); Sec 3) Techniques for evaluating factuality and its quantitative assessment (Huang et al., 2023b; Min et al., 2023); Sec 4) Analyzing the underlying mechanisms of factuality in LLMs and identifying the root causes of factual errors  (Liu et al., 2023b; Kotha et al., 2023); and Sec 5) Approaches to enhance the factuality of LLMs (Du et al., 2023; He et al., 2022). Notably, we categorize the use of LLMs into two primary settings: LLMs without external knowledge, such as ChatGPT (OpenAI, 2022b) and Retrieval-Augmented LLMs, such as BingChat (Microsoft, 2023). The complete structure of this survey is illustrated in Figure 1. Through a detailed examination of existing research, we seek to shed light on this critical aspect of LLMs, helping researchers, developers, and users harness the power of these models responsibly and effectively.
2 Factuality Issue
@@
2.1 Large Language Models
@@
  • General Domain LLMs: GPT-2 (Radford et al., 2019), GPT-3 (Brown et al., 2020), ChatGPT (OpenAI, 2022b), GPT-4 (OpenAI, 2023), GPT-Neo (Black et al., 2021), OPT (Zhang et al., 2022b), LLaMA (Touvron et al., 2023a), LLaMA-2 (Touvron et al., 2023b), Incite (Computer, 2023; Together, 2023), Falcon (Almazrouei et al., 2023), MPT Team (2023b), Vicuna (Chiang et al., 2023), FLAN-T5 Chung et al. (2022), BLOOM (Scao et al., 2022), Baichuan & Baichuan2 (Yang et al., 2023a), PaLM (Chowdhery et al., 2022), Gopher (Rae et al., 2022a), Megatron-LM (Shoeybi et al., 2019), SAIL (Luo et al., 2023a), Codex (Chen et al., 2021a), Bard (Google, 2023), GLM & ChatGLM (Zeng et al., 2022), InternLM (Team, 2023a), StableBeluga (Mahan et al., ), Claude (Ant, ), Alpaca (Taori et al., 2023), New Bing (Microsoft, 2023), Ziya-LLaMA (Zhang et al., 2022a), BLOOMZ (Muennighoff et al., 2022), Chinese-LLaMA (Cui et al., 2023c), Phoenix (Chen et al., 2023d), and others.
@@
  • Domain-specify LLMs: BloombergGPT (Wu et al., 2023), EcomGPT (Li et al., 2023e), BioGPT (Luo et al., 2022), LawGPT (Nguyen, 2023), Lawyer LLaMA (Huang et al., 2023a), ChatLaw (Cui et al., 2023b), BioMedLM (Venigalla et al., 2022), HuatuoGPT (Zhang et al., 2023a), ChatDoctor (Li et al., 2023f), MedicalGPT (Xu, 2023), Bentsao (Huatuo as its original name) (Wang et al., 2023c), Zhongjing (Yang et al., 2023c), LLM-AMT (Wang et al., 2023f), DISC-MedLLM (Bao et al., 2023), Cohortgpt  (Guan et al., 2023), Deid-gpt (Liu et al., 2023e), Doctorglm (Xiong et al., 2023b), MedChatZH (Tan et al., 2023a), K2 (Deng et al., 2023), HouYi (Bai et al., 2023), GrammarGPT (Fan et al., 2023), FoodGPT (Qi et al., 2023), ChatHome (Wen et al., 2023), and others.
2.2 Factuality
@@
TABLE I: Examples of different kinds of factual error produced by large language models. We category the factual error types by the causes of them, whose details can be found in Sec 4.2
@@
. Category Cause Example Dialog Notes and references Domain knowledge deficit Q: CEO of Assicurazioni Generali? BloombergGPT: Philippe Donnet GPT-NeoX: Antonio De Lorenzo, Simone Gambarini, Enrico Zanetti FLAN-T5-XXL: John M Forsyth, Christopher K Peters, {empty string} BloombergGPT is a finance domain-specific language model. Wu et al. (2023) Outdated information Q: When was Kyiv attacked by Russia? ChatGPT:As of my last knowledge update in September 2021, Russia had not launched an attack on Kyiv. Kyiv was attacked by Russia on 25 February 2022. Model- level causes Reasoning error Q: Who is Tom Cruise’s mother? A: Mary Lee Pfeiffer Q: Who is Mary Lee Pfeiffer’s son? A: There is no widely known information about… From Berglund et al. (2023). It is clear that the model knows Tom Cruise’s mother is Lee Pfeiffer, but it fails to reason that Lee Pfeiffer has a son named Tom Cruise. Distracted by the retrieval information Retrieved context: “Alan Melikdjanian - Age, Bio, Faces and Birthday” Melikdjanian was the Director, Writer (with his father), Editor, and Composer of the 2006 direct-to-DVD Russian-language film Citizen Mavzik, the story of a Russian immigrant family’s assimilation into American culture. The film was produced by Vilalan Productions, named after Melikdjanian and his father Vilen. Q: What is the nationality of the director of Citizen Mavzik? A: (1) The director of the film Citizen Mavzik is Alan Melikdjanian. (2) Alan Melikdjanian is from Russia. So, the answer is Russia. Alan Melikdjanian, known by the alias Captain Disillusion, is a Soviet-born American independent filmmaker and YouTuber (Melikdjanian, 2018). Misunderstood the retrieval information Retrieved context: “Gerard III, Count of Guelders - Wikipedia” Gerard married Margaretha of Brabant, the daughter of Duke Hendrik I of Brabant and Matilda of Flanders, Duchess of Brabant, in 1206. Gerard and Margaret’s Tomb … Q: Who is Gerard I, Count Of Guelders’s wife? A: … the name of Gerard I, Count of Guelders’s wife, which is Margaretha of Brabant. The retrieved context accurately provided the name of Gerard III, Count of Guelders’ wife, which is Margaretha of Brabant. However, the model erroneously referred to him as Gerard I, resulting in an incorrect answer. Retrieval- level causes Fail to address misinformation in retrieved documents Retrieved context: [0] “James Vincent - Death Notice and Service Information - Legacy.com” James Vincent died on August 9, 2023 in Long Beach, California Obituary. James Joseph Vincent was the perfect husband, father, grandfather,… [1] “UPDATE: Jury convicts James Vincent of murder after deliberating ” A jury convicted James Vincent of murder in the 2018 shooting death of Luka Grabb. The jury reached its verdict at 4:30 p.m. Friday. Q: When did James Vincent die? A: Based on the information provided in the two articles, James Vincent died on August 9, 2023, in Long Beach, California. In the presence of contradictory information from retrieved documents, the model should thoroughly outline the response, while also noting the need for clarity. For instance, the model could state: “According to the first document, James Vincent passed away on August 9, 2023, in Long Beach. However, the second document suggests that James Vincent was murdered in 2018.” Snowballing Q: Was there ever a US senator who represented the state of New Hampshire and whose alma mater was the University of Pennsylvania? A: Yes… His name was John P. Hale John P. Hale was graduated from Bowdoin College. Zhang et al. (2023b) Inference- level causes Exposure bias Q: Aamari was very stoic. [PRONOUN] rarely showed any emotion. A: He. The correct answer was Xe according to Hossain et al. (2023).
@@
The factuality problem refers to the potential of large language models to generate content that contradicts factual information. Here, “factual information” encompasses world knowledge, commonsense, and explicit information from reliable sources, such as Wikipedia or textbooks from specific domains  [^1] We only consider cases where the meaning is clear and the truthfulness can be determined in this survey. Furthermore, we only take into account undisputed facts. Minor errors that may exist in reliable sources are not within the scope of our consideration. . We showcase various instances of factuality issues in LLMs in Table I. For instance, an LLM might be deficient in domain-specific factual knowledge, such as medicine or law domain. Additionally, the LLM might be unaware of facts that occurred post its last update. There are also instances where the LLM, despite possessing the relevant facts, fails to reason out the correct answer. In some cases, it might even forget or be unable to recall facts it has previously learned.
@@
The factuality problem is closely related to several hot topics in the field of Large Language Models, including Hallucinations (Ji et al., 2023a; Zhao et al., 2023b), Outdated Information (Nakano et al., 2022; Qin et al., 2023), and Domain-Specificity (e.g., Health (Xiong et al., 2023a; Wang et al., 2023c), Law (Cui et al., 2023b), Finance (Wu et al., 2023)). At their core, these topics address the same issue: the potential for LLMs to generate contents that contradicts certain facts, whether those contents arise out of thin air, outdated information, or a lack of domain-specific knowledge. Therefore, we consider these three topics to fall within the scope of the factuality problem.
TABLE II: Comparison between the factuality issue and the hallucination issue.
Factual and Non-Hallucinated Factually correct outputs.
Non-Factual and Hallucinated Entirely fabricated outputs.
Hallucinated but Factual
1. Outputs that are unfaithful to the prompt but remain factually correct (Cao et al., 2022).
2. Outputs that deviate from the prompt’s specifics but don’t touch on factuality, e.g.,
a prompt asking for a story about a rabbit and wolf becoming friends, but the LLM
produces a tale about a rabbit and a dog befriending each other.
3. Outputs that provide additional factual details not specified in the prompt, e.g.,
a prompt asking about the capital of France, and the LLM responds with ”Paris, which
is known for the Eiffel Tower.”
Non-Factual but Non-Hallucinated
1. Outputs where the LLM states“I don’t know” or avoids a direct answer.
2. Outputs that are partially correct, e.g., for the question, ”Who landed on the moon
with Apollo 11?” If the LLM responds with just ”Neil Armstrong,” the answer is
incomplete but not hallucinated.
3. Outputs that provide a generalized or vague response without specific details, e.g.,
for a question about the causes of World War II, the LLM might respond with “It was
due to various political and economic factors.”
However, it is important to note that while these topics are related, they each have a unique focus.

### v3

@@ Abstract.
Factuality, factuality in LLM, retrievel augmented LLM, domain factuality enhanced LLM
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: CSURccs: Computing methodologies Natural language generation
1. Introduction
@@
, leaf, text width=15em ] ] [ Retrieval Aug-
mented LLMs [ BingChat (Microsoft, 2023), Llama-Index (Liu, 2022) , leaf, text width=15em ] ] ] [ Impact (Sec 2.3) [ Legal: Curran et al. (2023), Pranshu Verma (2023), Sands (2023) , leaf, text width=24em ] [ Medicine: Nori et al. (2023), Shen et al. (2023), Kung et al. (2023) , leaf, text width=24em ] ] ] [ Evaluation
@@
Metrics 3.1 [ Rule-based evaluation metrics; Neural evaluation metrics;
Human evaluation metrics; LLM-based evaluation metrics, leaf, text width=24em ] ] [ Factuality Evaluation
@@
C-Eval (Huang et al., 2023a), AGIEval (Zhong et al., 2023a),
RealTimeQA (Kasai et al., 2022), FreshQA (Vu et al., 2023), leaf, text width=24em ] ] [ Factuality
@@
Wang et al. (2023b), Yu et al. (2023), Kadavath et al. (2022), leaf, text width=24em ] ] [ Domain-Factuality
Evaluation (Sec 3.4) [ FLARE (Finance) (Xie et al., 2023a), CMB (Medicine) (Wang et al., 2023a),
GeneTuring (Gene) (Jin et al., 2023), EcomInstruct (Ecom) (Li et al., 2023c), leaf, text width=24em ] ] ] [ Analysis
@@
Immemorization, Forgetting, Reasoning Failure, leaf, text width=18em ] ] [ Retrieval-Level 4.2.2 [ Insufficient Information, Distracting Information,
@@
Misinterpretation of Related Information, leaf, text width=18em ] ] [ Inference-Level 4.2.3 [ Snowballing, Erroneous Decoding, Exposure Bias, leaf, text width=18em ] ] ] ] [ Enhancement
@@
, leaf, text width=18em ] ] [ SFT-based 5.1.2 [ Continual SFT: Moiseev et al. (2022);
@@
, leaf, text width=18em ] ] [ Multi-Agent 5.1.3 [ Multi-debate: Du et al. (2023)
@@
Agent-based Retrieval: ReAct (Yao et al., 2023b) , leaf, text width=16em ] ] [ Retrieval
@@
, leaf, text width=16em ] ] [ Retrieval on External
Memory 5.2.4 [ General Knowledge: Li et al. (2022a),
@@
Entity Knowledge: knowledge plugin (Zhang et al., 2023h)
@@
, leaf, text width=16em ] ] [ Retrieval on
Structured Data 5.2.5 [ StructGPT (Jiang et al., 2023b) , Zhang et al. (2023f) , leaf, text width=16em ] ] ] [ Domain Factuality
@@
Finance: EcomGPT (Li et al., 2023c), BloombergGPT (Wu et al., 2023)
@@
, leaf, text width=18em ] ] ] ]
Figure 1. Taxonomy of research on factuality in Large Language Models that consists of the issue, evaluation, analysis and enhancement.
@@
Amidst these pioneering efforts, the emergence of Large Language Models (LLMs), such as ChatGPT (OpenAI, 2022b), GPT-4 (OpenAI, 2023) and LLaMA (Touvron et al., 2023a, b), has been seen as a significant leap in both academics and industries, especially towards AI systems possessing vast factual knowledge (Petroni et al., 2019a; De Cao et al., 2021; OpenAI, 2023). The advantages of using LLMs as knowledge bases carriers are manifold. Firstly, they reduce the overhead and costs associated with building and maintaining dedicated knowledge bases (Petroni et al., 2019c; AlKhamissi et al., 2022; Wang et al., 2023e). Additionally, LLMs offer a more flexible approach to knowledge processing and utilization, allowing for context-aware reasoning and the ability to adapt to novel information or prompts (Sun et al., 2023a; Huang and Chang, 2023). Yet, with their unparalleled capabilities, concerns have arisen about the potential of LLMs to generate non-factual or misleading content (Bender et al., 2021; OpenAI, 2023; Bubeck et al., 2023). In light of these advancements and challenges, this survey seeks to delve deeply into the LLMs, exploring both their potential and the concerns surrounding their factual accuracy.
@@
Understanding the factuality of Large Language Models is more than just a technical challenge; it’s essential for the responsible use of these tools in our daily lives. As LLMs become more integrated into services like search engines (Microsoft, 2023), chatbots (OpenAI, 2022b; Google, 2023), and content generators (Cui et al., 2023b), the information they provide directly influences decisions, beliefs, and actions of millions of people. If an LLM provides incorrect or misleading information, it can lead to misunderstandings, spread false beliefs, or even cause harm, especially for those domains that demand high factual accuracy (Ling et al., 2023b), such as health  (Thirunavukarasu et al., 2023; Tang et al., 2023), law (Huang et al., 2023c), and finance (Wu et al., 2023). For instance, a physician relying on an LLM for medical guidance might inadvertently jeopardize patient health, a corporation leveraging LLM insights might make ill-informed market decisions, or an attorney misinformed by an LLM might falter in legal proceedings (Curran et al., 2023). In addition, with the advancement of LLM-based agents, the factuality of LLMs is becoming even more potent. A driver or an autonomous driving car might rely on LLM-based agents for planning or driving, where serious factual mistakes made by LLMs could cause irreversible damage. By studying the factuality of LLMs, we aim to ensure that these models are both powerful and trustworthy.
@@
A surge of research has been directed towards evaluating LLMs’ factuality, which encompasses diverse tasks like factoid question answering and fact checking. Beyond evaluation, efforts to improve the factual knowledge of LLMs have been notable. Strategies have ranged from retrieving information from external knowledge bases to continual pretraining and supervised finetuning. Yet, despite these burgeoning efforts, a holistic overview that covers the full spectrum of factuality in LLMs remains elusive. While there are existing surveys in the field, such as those by Chang et al. (2023) and Wang et al. (2023i), that delve into the evaluation of LLMs and their factuality, they only scratch the surface of the broader landscape. There are also a bunch of recent studies focusing on hallucinations in LLMs (Rawte et al., 2023; Zhang et al., 2023c; Rawte et al., 2023; Ye et al., 2023; Huang et al., 2023d). But we differentiate between the hallucination issue and the factuality issues in Sec 2.2. Moreover, these surveys often overlook key areas we emphasize, like domain-specific factuality or the challenge of outdated information. While Ling et al. (2023a) explores domain specialization in LLMs, our survey takes a more expansive look at the broader issues of factuality. To the best of our understanding, our work is the first comprehensive study on the factuality of large language models.
@@
This survey aims to offer an exhaustive overview of the factuality studies in LLMs, delving into four key dimensions: Sec 2) The definition and impact of the factuality issue (Pranshu Verma, 2023; Nori et al., 2023); Sec 3) Techniques for evaluating factuality and its quantitative assessment (Huang et al., 2023a; Min et al., 2023); Sec 4) Analyzing the underlying mechanisms of factuality in LLMs and identifying the root causes of factual errors  (Liu et al., 2023b; Kotha et al., 2023); and Sec 5) Approaches to enhance the factuality of LLMs (Du et al., 2023; He et al., 2022). Notably, we categorize the use of LLMs into two primary settings: LLMs without external knowledge, such as ChatGPT (OpenAI, 2022b) and Retrieval-Augmented LLMs, such as BingChat (Microsoft, 2023). The complete structure of this survey is illustrated in Figure 1. Through a detailed examination of existing research, we seek to shed light on this critical aspect of LLMs, helping researchers, developers, and users harness the power of these models responsibly and effectively.
2. Factuality Issue
@@
2.1. Large Language Models
@@
  • General Domain LLMs: GPT-2 (Radford et al., 2019), GPT-3 (Brown et al., 2020), ChatGPT (OpenAI, 2022b), GPT-4 (OpenAI, 2023), GPT-Neo (Black et al., 2021), OPT (Zhang et al., 2022b), LLaMA (Touvron et al., 2023a), LLaMA-2 (Touvron et al., 2023b), Incite (Computer, 2023; Together, 2023), Claude (Claude, 2023), Falcon (Almazrouei et al., 2023), MPT (Team, 2023b), Vicuna (Chiang et al., 2023), FLAN-T5 (Chung et al., 2022), BLOOM (Scao et al., 2022), Baichuan & Baichuan2 (Yang et al., 2023b), PaLM (Chowdhery et al., 2022), Gopher (et al., [n. d.]), Megatron-LM (Shoeybi et al., 2019), SAIL (Luo et al., 2023a), Codex (Chen et al., 2021b), Bard (Google, 2023), GLM & ChatGLM (Zeng et al., 2022), InternLM (Team, 2023a), StableBeluga (Mahan et al., [n. d.]), Claude (Claude, 2023), Alpaca (Taori et al., 2023), New Bing (Microsoft, 2023), Ziya-LLaMA (Zhang et al., 2022a), BLOOMZ (Muennighoff et al., 2022), Chinese-LLaMA (Cui et al., 2023c), Phoenix (Chen et al., 2023c), and others.
@@
  • Domain-specify LLMs: BloombergGPT (Wu et al., 2023), EcomGPT (Li et al., 2023c), BioGPT (Luo et al., 2022), LawGPT (Nguyen, 2023), Lawyer LLaMA (Huang et al., 2023c), ChatLaw (Cui et al., 2023b), BioMedLM (Venigalla et al., 2022), HuatuoGPT (Zhang et al., 2023a), ChatDoctor (Li et al., 2023b), MedicalGPT (Xu, 2023), Bentsao (Huatuo as its original name) (Wang et al., 2023d), Zhongjing (Yang et al., 2023c), LLM-AMT (Wang et al., 2023f), DISC-MedLLM (Bao et al., 2023), Cohortgpt  (Guan et al., 2023), Deid-gpt (Liu et al., 2023e), Doctorglm (Xiong et al., 2023b), MedChatZH (Tan et al., 2023a), K2 (Deng et al., 2023), HouYi (Bai et al., 2023), GrammarGPT (Fan et al., 2023), FoodGPT (Qi et al., 2023), ChatHome (Wen et al., 2023), and others.
2.2. Factuality
@@
Table 1. Examples of different kinds of factual error produced by large language models. We category the factual error types by the causes of them, whose details can be found in Sec 4.2
@@
Category Cause Example Dialog Notes and references Domain knowledge deficit Q: CEO of Assicurazioni Generali? BloombergGPT: Philippe Donnet GPT-NeoX: Antonio De Lorenzo, Simone Gambarini, Enrico Zanetti FLAN-T5-XXL: John M Forsyth, Christopher K Peters, {empty string} BloombergGPT is a finance domain-specific language model. (Wu et al., 2023) Outdated information Q: When was Kyiv attacked by Russia? ChatGPT:As of my last knowledge update in September 2021, Russia had not launched an attack on Kyiv. Kyiv was attacked by Russia on 25 February 2022. Model- level causes Reasoning error Q: Who is Tom Cruise’s mother? A: Mary Lee Pfeiffer Q: Who is Mary Lee Pfeiffer’s son? A: There is no widely known information about… From Berglund et al. (2023). It is clear that the model knows Tom Cruise’s mother is Lee Pfeiffer, but it fails to reason that Lee Pfeiffer has a son named Tom Cruise. Distracted by the retrieval information Retrieved context: “Alan Melikdjanian - Age, Bio, Faces and Birthday” Melikdjanian was the Director, Writer (with his father), Editor, and Composer of the 2006 direct-to-DVD Russian-language film Citizen Mavzik, the story of a Russian immigrant family’s assimilation into American culture. The film was produced by Vilalan Productions, named after Melikdjanian and his father Vilen. Q: What is the nationality of the director of Citizen Mavzik? A: (1) The director of the film Citizen Mavzik is Alan Melikdjanian. (2) Alan Melikdjanian is from Russia. So, the answer is Russia. Alan Melikdjanian, known by the alias Captain Disillusion, is a Soviet-born American independent filmmaker and YouTuber (Melikdjanian, 2018). Misunderstood the retrieval information Retrieved context: “Gerard III, Count of Guelders - Wikipedia” Gerard married Margaretha of Brabant, the daughter of Duke Hendrik I of Brabant and Matilda of Flanders, Duchess of Brabant, in 1206. Gerard and Margaret’s Tomb … Q: Who is Gerard I, Count Of Guelders’s wife? A: … the name of Gerard I, Count of Guelders’s wife, which is Margaretha of Brabant. The retrieved context accurately provided the name of Gerard III, Count of Guelders’ wife, which is Margaretha of Brabant. However, the model erroneously referred to him as Gerard I, resulting in an incorrect answer. Retrieval- level causes Fail to address misinformation in retrieved documents Retrieved context: [0] “James Vincent - Death Notice and Service Information - Legacy.com” James Vincent died on August 9, 2023 in Long Beach, California Obituary. James Joseph Vincent was the perfect husband, father, grandfather,… [1] “UPDATE: Jury convicts James Vincent of murder after deliberating ” A jury convicted James Vincent of murder in the 2018 shooting death of Luka Grabb. The jury reached its verdict at 4:30 p.m. Friday. Q: When did James Vincent die? A: Based on the information provided in the two articles, James Vincent died on August 9, 2023, in Long Beach, California. In the presence of contradictory information from retrieved documents, the model should thoroughly outline the response, while also noting the need for clarity. For instance, the model could state: “According to the first document, James Vincent passed away on August 9, 2023, in Long Beach. However, the second document suggests that James Vincent was murdered in 2018.” Snowballing Q: Was there ever a US senator who represented the state of New Hampshire and whose alma mater was the University of Pennsylvania? A: Yes… His name was John P. Hale John P. Hale was graduated from Bowdoin College. (Zhang et al., 2023g) Inference- level causes Exposure bias Q: Aamari was very stoic. [PRONOUN] rarely showed any emotion. A: He. The correct answer was Xe according to (Hossain et al., 2023).
@@
By factuality in LLMs we refer to the capability of large language models for generating contents that follow factual information, which encompasses commonsense, world knowledge and domain facts. The factual information can be grounded to reliable sources, such as dictionaries, Wikipedia or textbooks from different domains  [^1] We only consider cases where the meaning is clear and the truthfulness can be determined in this survey. Furthermore, we only take into account undisputed facts. Minor errors that may exist in reliable sources are not within the scope of our consideration. . A series of work have discussed whether LLMs can serve as knowledge bases to store factual knowledge (AlKhamissi et al., 2022; Yu et al., 2023; Pan et al., 2023a)
Table 2. Comparison between the factuality issue and the hallucination issue.
Factual and Non-Hallucinated Factually correct outputs. Non-Factual and Hallucinated Entirely fabricated outputs. Hallucinated but Factual 1. Outputs that are unfaithful to the prompt but remain factually correct (Cao et al., 2022). 2. Outputs that deviate from the prompt’s specifics but don’t touch on factuality, e.g., a prompt asking for a story about a rabbit and wolf becoming friends, but the LLM produces a tale about a rabbit and a dog befriending each other. 3. Outputs that provide additional factual details not specified in the prompt, e.g., a prompt asking about the capital of France, and the LLM responds with ”Paris, which is known for the Eiffel Tower.” Non-Factual but Non-Hallucinated 1. Outputs where the LLM states“I don’t know” or avoids a direct answer. 2. Outputs that are partially correct, e.g., for the question, ”Who landed on the moon with Apollo 11?” If the LLM responds with just ”Neil Armstrong,” the answer is incomplete but not hallucinated. 3. Outputs that provide a generalized or vague response without specific details, e.g., for a question about the causes of World War II, the LLM might respond with “It was due to various political and economic factors.”
@@
Existing work focus on measuring factuality in LLMs qualitatively (Lin et al., 2022b; Chern et al., 2023), discussing the mechanism for storing knowledge (Meng et al., 2022; Chen et al., 2023a) and tracing the source of knowledge issues (Gou et al., 2023; Kandpal et al., 2023). The factuality issue for LLMs receive relatively the most attention. Several instances are shown in Table 1. For instance, an LLM might be deficient in domain-specific factual knowledge, such as medicine or law domain. Additionally, the LLM might be unaware of facts that occurred post its last update. There are also instances where the LLM, despite possessing the relevant facts, fails to reason out the correct answer. In some cases, it might even forget or be unable to recall facts it has previously learned. The factuality problem is closely related to several hot topics in the field of Large Language Models, including Hallucinations (Ji et al., 2023a; Zhao et al., 2023b), Outdated Information (Nakano et al., 2022; Qin et al., 2023), and Domain-Specificity (e.g., Health (Xiong et al., 2023a; Wang et al., 2023d), Law (Cui et al., 2023b), Finance (Wu et al., 2023)). At their core, these topics address the same issue: the potential for LLMs to generate contents that contradicts certain facts, whether those contents arise out of thin air, outdated information, or a lack of domain-specific knowledge.